OpenStudy (anonymous):

What is the kinetic energy of a toy truck with a mass of 0.75 kg and a velocity of 4 m/s? (Formula: ) 3 J 6 J 12 J 24 J

OpenStudy (anonymous):

K.E = 1/2 m v square so K.E = 1/2 0.75 kg * 4 m/s K.E = 1.50 kg m/s K.E = 1.5 N